OpenStudy (anonymous):

Find the center and vertices of the hyperbola. 16x^2 - 9y^2 - 96x + 72y - 144 = 0

OpenStudy (amistre64):

\[(\frac{x}{a})^2-(\frac{y}{b})^2=1\]